SOLVE THE SYSTEM BY ADDITION
X+Y=6
X-Y=4
THANKS FOR ALL YOUR HELP

Answer by askmemath(368)

X+Y=6
X-Y=4
Adding the 2 equations
2X = 10
Dividing both sides by 2
X = 5
=> Y = 1 since 5+Y = 6
=>y = 6-5
=>Y = 1